\begin{thebibliography}{55}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Althnian et~al.(2021)Althnian, AlSaeed, Al-Baity, Samha, Dris, Alzakari, Abou~Elwafa, and Kurdi}]{althnian2021impact}
Alhanoof Althnian, Duaa AlSaeed, Heyam Al-Baity, Amani Samha, Alanoud~Bin Dris, Najla Alzakari, Afnan Abou~Elwafa, and Heba Kurdi. 2021.
\newblock Impact of dataset size on classification performance: an empirical evaluation in the medical domain.
\newblock \emph{Applied Sciences}, 11(2):796.

\bibitem[{Anaby-Tavor et~al.(2020)Anaby-Tavor, Carmeli, Goldbraich, Kantor, Kour, Shlomov, Tepper, and Zwerdling}]{anaby2020not}
Ateret Anaby-Tavor, Boaz Carmeli, Esther Goldbraich, Amir Kantor, George Kour, Segev Shlomov, Naama Tepper, and Naama Zwerdling. 2020.
\newblock Do not have enough data? deep learning to the rescue!
\newblock In \emph{Proceedings of the AAAI Conference on Artificial Intelligence}, volume~34, pages 7383--7390.

\bibitem[{Bayer et~al.(2022)Bayer, Kaufhold, and Reuter}]{bayer2022survey}
Markus Bayer, Marc-Andr{\'e} Kaufhold, and Christian Reuter. 2022.
\newblock A survey on data augmentation for text classification.
\newblock \emph{ACM Computing Surveys}, 55(7):1--39.

\bibitem[{Belinkov and Bisk(2018)}]{belinkov2018synthetic}
Yonatan Belinkov and Yonatan Bisk. 2018.
\newblock Synthetic and natural noise both break neural machine translation.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Bergstra et~al.(2011)Bergstra, Bardenet, Bengio, and K{\'e}gl}]{bergstra2011algorithms}
James Bergstra, R{\'e}mi Bardenet, Yoshua Bengio, and Bal{\'a}zs K{\'e}gl. 2011.
\newblock Algorithms for hyper-parameter optimization.
\newblock \emph{Advances in neural information processing systems}, 24.

\bibitem[{Bergstra and Bengio(2012)}]{bergstra2012random}
James Bergstra and Yoshua Bengio. 2012.
\newblock Random search for hyper-parameter optimization.
\newblock \emph{Journal of machine learning research}, 13(2).

\bibitem[{Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell et~al.}]{brown2020language}
Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared~D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et~al. 2020.
\newblock Language models are few-shot learners.
\newblock \emph{Advances in neural information processing systems}, 33:1877--1901.

\bibitem[{Chen et~al.(2023)Chen, Tam, Raffel, Bansal, and Yang}]{chen2023empirical}
Jiaao Chen, Derek Tam, Colin Raffel, Mohit Bansal, and Diyi Yang. 2023.
\newblock An empirical survey of data augmentation for limited data learning in nlp.
\newblock \emph{Transactions of the Association for Computational Linguistics}, 11:191--211.

\bibitem[{Choi et~al.(2023)Choi, Jin, Lee, Song, and Kim}]{choi2023softeda}
Juhwan Choi, Kyohoon Jin, Junho Lee, Sangmin Song, and YoungBin Kim. 2023.
\newblock Softeda: Rethinking rule-based data augmentation with soft labels.
\newblock In \emph{ICLR 2023 Tiny Papers}.

\bibitem[{Cubuk et~al.(2019)Cubuk, Zoph, Mane, Vasudevan, and Le}]{cubuk2019autoaugment}
Ekin~D Cubuk, Barret Zoph, Dandelion Mane, Vijay Vasudevan, and Quoc~V Le. 2019.
\newblock Autoaugment: Learning augmentation strategies from data.
\newblock In \emph{Proceedings of the IEEE/CVF conference on computer vision and pattern recognition}, pages 113--123.

\bibitem[{Dai et~al.(2023)Dai, Liu, Liao, Huang, Cao, Wu, Zhao, Xu, Liu, Liu et~al.}]{dai2023auggpt}
Haixing Dai, Zhengliang Liu, Wenxiong Liao, Xiaoke Huang, Yihan Cao, Zihao Wu, Lin Zhao, Shaochen Xu, Wei Liu, Ninghao Liu, et~al. 2023.
\newblock Auggpt: Leveraging chatgpt for text data augmentation.
\newblock \emph{arXiv preprint arXiv:2302.13007}.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and Toutanova}]{devlin2019bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock Bert: Pre-training of deep bidirectional transformers for language understanding.
\newblock In \emph{Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers)}, pages 4171--4186.

\bibitem[{Feng et~al.(2021)Feng, Gangal, Wei, Chandar, Vosoughi, Mitamura, and Hovy}]{feng2021survey}
Steven~Y Feng, Varun Gangal, Jason Wei, Sarath Chandar, Soroush Vosoughi, Teruko Mitamura, and Eduard Hovy. 2021.
\newblock A survey of data augmentation approaches for nlp.
\newblock In \emph{Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021}, pages 968--988.

\bibitem[{Ganapathibhotla and Liu(2008)}]{ganapathibhotla-liu-2008-mining}
Murthy Ganapathibhotla and Bing Liu. 2008.
\newblock Mining opinions in comparative sentences.
\newblock In \emph{Proceedings of the 22nd International Conference on Computational Linguistics (Coling 2008)}, pages 241--248.

\bibitem[{Guo(2020)}]{guo2020nonlinear}
Hongyu Guo. 2020.
\newblock Nonlinear mixup: Out-of-manifold data augmentation for text classification.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial Intelligence}, volume~34, pages 4044--4051.

\bibitem[{He et~al.(2023)He, Gao, and Chen}]{he2022debertav3}
Pengcheng He, Jianfeng Gao, and Weizhu Chen. 2023.
\newblock Debertav3: Improving deberta using electra-style pre-training with gradient-disentangled embedding sharing.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{He et~al.(2021)He, Liu, Gao, and Chen}]{he2020deberta}
Pengcheng He, Xiaodong Liu, Jianfeng Gao, and Weizhu Chen. 2021.
\newblock Deberta: Decoding-enhanced bert with disentangled attention.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Hern{\'a}ndez-Garc{\'\i}a and K{\"o}nig(2018)}]{hernandez2018data}
Alex Hern{\'a}ndez-Garc{\'\i}a and Peter K{\"o}nig. 2018.
\newblock Data augmentation instead of explicit regularization.
\newblock \emph{arXiv preprint arXiv:1806.03852}.

\bibitem[{Hu and Liu(2004)}]{hu2004mining}
Minqing Hu and Bing Liu. 2004.
\newblock Mining and summarizing customer reviews.
\newblock In \emph{Proceedings of the Tenth ACM SIGKDD International Conference on Knowledge Discovery and Data Mining}, page 168–177.

\bibitem[{Hu et~al.(2019)Hu, Tan, Salakhutdinov, Mitchell, and Xing}]{hu2019learning}
Zhiting Hu, Bowen Tan, Russ~R Salakhutdinov, Tom~M Mitchell, and Eric~P Xing. 2019.
\newblock Learning data manipulation for augmentation and weighting.
\newblock \emph{Advances in Neural Information Processing Systems}, 32.

\bibitem[{Karimi et~al.(2021)Karimi, Rossi, and Prati}]{karimi2021aeda}
Akbar Karimi, Leonardo Rossi, and Andrea Prati. 2021.
\newblock Aeda: An easier data augmentation technique for text classification.
\newblock In \emph{Findings of the Association for Computational Linguistics: EMNLP 2021}, pages 2748--2754.

\bibitem[{Kim et~al.(2022)Kim, Kang, Ahn, and Shin}]{kim2021makes}
Jaehyung Kim, Dongyeop Kang, Sungsoo Ahn, and Jinwoo Shin. 2022.
\newblock What makes better augmentation strategies? augment difficult but not too different.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Kim et~al.(2021)Kim, Jeong, and Cho}]{kim2021linda}
Yekyung Kim, Seohyeong Jeong, and Kyunghyun Cho. 2021.
\newblock Linda: Unsupervised learning to interpolate in natural language processing.
\newblock \emph{arXiv preprint arXiv:2112.13969}.

\bibitem[{Li et~al.(2022)Li, Hou, and Che}]{li2022data}
Bohan Li, Yutai Hou, and Wanxiang Che. 2022.
\newblock Data augmentation approaches in natural language processing: A survey.
\newblock \emph{AI Open}, 3:71--90.

\bibitem[{Li and Roth(2002)}]{li2002learning}
Xin Li and Dan Roth. 2002.
\newblock Learning question classifiers.
\newblock In \emph{COLING 2002: The 19th International Conference on Computational Linguistics}.

\bibitem[{Liaw et~al.(2018)Liaw, Liang, Nishihara, Moritz, Gonzalez, and Stoica}]{liaw2018tune}
Richard Liaw, Eric Liang, Robert Nishihara, Philipp Moritz, Joseph~E Gonzalez, and Ion Stoica. 2018.
\newblock Tune: A research platform for distributed model selection and training.
\newblock \emph{arXiv preprint arXiv:1807.05118}.

\bibitem[{Liu et~al.(2015)Liu, Gao, Liu, and Zhang}]{qian2015automated}
Qian Liu, Zhiqiang Gao, Bing Liu, and Yuanlin Zhang. 2015.
\newblock Automated rule selection for aspect extraction in opinion mining.
\newblock In \emph{Proceedings of the 24th International Conference on Artificial Intelligence}, page 1291–1297.

\bibitem[{Longpre et~al.(2020)Longpre, Wang, and DuBois}]{longpre2020effective}
Shayne Longpre, Yu~Wang, and Chris DuBois. 2020.
\newblock How effective is task-agnostic data augmentation for pretrained transformers?
\newblock In \emph{Findings of the Association for Computational Linguistics: EMNLP 2020}, pages 4401--4411.

\bibitem[{Marivate and Sefara(2020)}]{marivate2020improving}
Vukosi Marivate and Tshephisho Sefara. 2020.
\newblock Improving short text classification through global augmentation methods.
\newblock In \emph{International Cross-Domain Conference for Machine Learning and Knowledge Extraction}, pages 385--399. Springer.

\bibitem[{Okimura et~al.(2022)Okimura, Reid, Kawano, and Matsuo}]{okimura2022impact}
Itsuki Okimura, Machel Reid, Makoto Kawano, and Yutaka Matsuo. 2022.
\newblock On the impact of data augmentation on downstream performance in natural language processing.
\newblock In \emph{Proceedings of the Third Workshop on Insights from Negative Results in NLP}, pages 88--93.

\bibitem[{Pang and Lee(2004)}]{pang2004sentimental}
Bo~Pang and Lillian Lee. 2004.
\newblock A sentimental education: Sentiment analysis using subjectivity summarization based on minimum cuts.
\newblock In \emph{Proceedings of the 42nd Annual Meeting of the Association for Computational Linguistics (ACL-04)}, pages 271--278.

\bibitem[{Pang et~al.(2002)Pang, Lee, and Vaithyanathan}]{pang-etal-2002-thumbs}
Bo~Pang, Lillian Lee, and Shivakumar Vaithyanathan. 2002.
\newblock Thumbs up? sentiment classification using machine learning techniques.
\newblock In \emph{Proceedings of the 2002 Conference on Empirical Methods in Natural Language Processing ({EMNLP} 2002)}, pages 79--86.

\bibitem[{Paszke et~al.(2019)Paszke, Gross, Massa, Lerer, Bradbury, Chanan, Killeen, Lin, Gimelshein, Antiga et~al.}]{paszke2019pytorch}
Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, et~al. 2019.
\newblock Pytorch: An imperative style, high-performance deep learning library.
\newblock \emph{Advances in neural information processing systems}, 32.

\bibitem[{Plu{\v{s}}{\v{c}}ec and {\v{S}}najder(2023)}]{pluvsvcec2023data}
Domagoj Plu{\v{s}}{\v{c}}ec and Jan {\v{S}}najder. 2023.
\newblock Data augmentation for neural nlp.
\newblock \emph{arXiv preprint arXiv:2302.11412}.

\bibitem[{Prusa et~al.(2015)Prusa, Khoshgoftaar, and Seliya}]{prusa2015effect}
Joseph Prusa, Taghi~M Khoshgoftaar, and Naeem Seliya. 2015.
\newblock The effect of dataset size on training tweet sentiment classifiers.
\newblock In \emph{2015 IEEE 14th International Conference on Machine Learning and Applications (ICMLA)}, pages 96--102. IEEE.

\bibitem[{Radford et~al.(2019)Radford, Wu, Child, Luan, Amodei, Sutskever et~al.}]{radford2019language}
Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et~al. 2019.
\newblock Language models are unsupervised multitask learners.
\newblock \emph{OpenAI blog}, 1(8):9.

\bibitem[{Raffel et~al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena, Zhou, Li, and Liu}]{raffel2020exploring}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter~J Liu. 2020.
\newblock Exploring the limits of transfer learning with a unified text-to-text transformer.
\newblock \emph{The Journal of Machine Learning Research}, 21(1):5485--5551.

\bibitem[{Ren et~al.(2021)Ren, Zhang, Li, Sun, and Zhou}]{ren2021text}
Shuhuai Ren, Jinchao Zhang, Lei Li, Xu~Sun, and Jie Zhou. 2021.
\newblock Text autoaugment: Learning compositional augmentation policy for text classification.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing}, pages 9029--9043.

\bibitem[{Sennrich et~al.(2016)Sennrich, Haddow, and Birch}]{sennrich2016improving}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016.
\newblock Improving neural machine translation models with monolingual data.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 86--96.

\bibitem[{Socher et~al.(2013)Socher, Perelygin, Wu, Chuang, Manning, Ng, and Potts}]{socher2013recursive}
Richard Socher, Alex Perelygin, Jean Wu, Jason Chuang, Christopher~D Manning, Andrew~Y Ng, and Christopher Potts. 2013.
\newblock Recursive deep models for semantic compositionality over a sentiment treebank.
\newblock In \emph{Proceedings of the 2013 conference on empirical methods in natural language processing}, pages 1631--1642.

\bibitem[{Sun et~al.(2020)Sun, Xia, Yin, Liang, Yu, and He}]{sun-etal-2020-mixup}
Lichao Sun, Congying Xia, Wenpeng Yin, Tingting Liang, Philip Yu, and Lifang He. 2020.
\newblock Mixup-transformer: Dynamic data augmentation for {NLP} tasks.
\newblock In \emph{Proceedings of the 28th International Conference on Computational Linguistics}, pages 3436--3440.

\bibitem[{Szegedy et~al.(2016)Szegedy, Vanhoucke, Ioffe, Shlens, and Wojna}]{szegedy2016rethinking}
Christian Szegedy, Vincent Vanhoucke, Sergey Ioffe, Jon Shlens, and Zbigniew Wojna. 2016.
\newblock Rethinking the inception architecture for computer vision.
\newblock In \emph{Proceedings of the IEEE conference on computer vision and pattern recognition}, pages 2818--2826.

\bibitem[{Warstadt et~al.(2019)Warstadt, Singh, and Bowman}]{warstadt2019neural}
Alex Warstadt, Amanpreet Singh, and Samuel~R Bowman. 2019.
\newblock Neural network acceptability judgments.
\newblock \emph{Transactions of the Association for Computational Linguistics}, 7:625--641.

\bibitem[{Wei and Zou(2019)}]{wei2019eda}
Jason Wei and Kai Zou. 2019.
\newblock Eda: Easy data augmentation techniques for boosting performance on text classification tasks.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP)}, pages 6382--6388.

\bibitem[{Wolf et~al.(2020)Wolf, Debut, Sanh, Chaumond, Delangue, Moi, Cistac, Rault, Louf, Funtowicz et~al.}]{wolf2020transformers}
Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, R{\'e}mi Louf, Morgan Funtowicz, et~al. 2020.
\newblock Transformers: State-of-the-art natural language processing.
\newblock In \emph{Proceedings of the 2020 conference on empirical methods in natural language processing: system demonstrations}, pages 38--45.

\bibitem[{Wu et~al.(2019)Wu, Lv, Zang, Han, and Hu}]{wu2019conditional}
Xing Wu, Shangwen Lv, Liangjun Zang, Jizhong Han, and Songlin Hu. 2019.
\newblock Conditional bert contextual augmentation.
\newblock In \emph{Computational Science--ICCS 2019: 19th International Conference, Faro, Portugal, June 12--14, 2019, Proceedings, Part IV 19}, pages 84--95. Springer.

\bibitem[{Yang et~al.(2022)Yang, Xiao, Zhang, Guo, Zhao, and Shen}]{yang2022image}
Suorong Yang, Weikang Xiao, Mengcheng Zhang, Suhan Guo, Jian Zhao, and Furao Shen. 2022.
\newblock Image data augmentation for deep learning: A survey.
\newblock \emph{arXiv preprint arXiv:2204.08610}.

\bibitem[{Yoo et~al.(2021)Yoo, Park, Kang, Lee, and Park}]{yoo2021gpt3mix}
Kang~Min Yoo, Dongju Park, Jaewook Kang, Sang-Woo Lee, and Woomyoung Park. 2021.
\newblock Gpt3mix: Leveraging large-scale language models for text augmentation.
\newblock In \emph{Findings of the Association for Computational Linguistics: EMNLP 2021}, pages 2225--2239.

\bibitem[{Yoon et~al.(2021)Yoon, Kim, and Park}]{yoon2021ssmix}
Soyoung Yoon, Gyuwan Kim, and Kyumin Park. 2021.
\newblock Ssmix: Saliency-based span mixup for text classification.
\newblock In \emph{Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021}, pages 3225--3234.

\bibitem[{Zhang et~al.(2018)Zhang, Cisse, Dauphin, and Lopez-Paz}]{zhang2018mixup}
Hongyi Zhang, Moustapha Cisse, Yann~N. Dauphin, and David Lopez-Paz. 2018.
\newblock mixup: Beyond empirical risk minimization.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Zhang et~al.(2022)Zhang, Yang, and Yang}]{zhang2022treemix}
Le~Zhang, Zichao Yang, and Diyi Yang. 2022.
\newblock Treemix: Compositional constituency-based data augmentation for natural language understanding.
\newblock In \emph{Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies}, pages 5243--5258.

\bibitem[{Zhang and Ma(2022)}]{zhang2022good}
Linfeng Zhang and Kaisheng Ma. 2022.
\newblock A good data augmentation policy is not all you need: A multi-task learning perspective.
\newblock \emph{IEEE Transactions on Circuits and Systems for Video Technology}.

\bibitem[{Zhang et~al.(2015)Zhang, Zhao, and LeCun}]{zhang2015character}
Xiang Zhang, Junbo Zhao, and Yann LeCun. 2015.
\newblock Character-level convolutional networks for text classification.
\newblock \emph{Advances in neural information processing systems}, 28.

\bibitem[{Zhao et~al.(2022)Zhao, Zhang, Xu, Ding, Guan, and Zhou}]{zhao2022epida}
Minyi Zhao, Lu~Zhang, Yi~Xu, Jiandong Ding, Jihong Guan, and Shuigeng Zhou. 2022.
\newblock Epida: An easy plug-in data augmentation framework for high performance text classification.
\newblock In \emph{Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies}, pages 4742--4752.

\bibitem[{Zhou et~al.(2022)Zhou, Zheng, Tang, Jian, and Yang}]{zhou2022flipda}
Jing Zhou, Yanan Zheng, Jie Tang, Li~Jian, and Zhilin Yang. 2022.
\newblock Flipda: Effective and robust data augmentation for few-shot learning.
\newblock In \emph{Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 8646--8665.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[preprint]{acl}

\usepackage{times}
\usepackage{latexsym}
\usepackage{makecell}
\usepackage{graphicx}
\usepackage{hyperref}
\usepackage{multirow} 
\usepackage{float} 
\usepackage{adjustbox} 
\usepackage{verbatim} 
\usepackage{amsmath} 
\usepackage[]{xcolor}
\usepackage{bm} 

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\title{AutoAugment Is What You Need: \\ Enhancing Rule-based Augmentation Methods in Low-resource Regimes}

\author{Juhwan Choi\textsuperscript{1}, Kyohoon Jin\textsuperscript{2}, Junho Lee\textsuperscript{1}, Sangmin Song\textsuperscript{1} \and Youngbin Kim\textsuperscript{1,2} \\ \\
  \textsuperscript{1}Department of Artificial Intelligence, Chung-Ang University \\
  \textsuperscript{2}Graduate School of Advanced Imaging Sciences, Multimedia and Film, Chung-Ang University \\
  \texttt{\{gold5230,fhzh123,jhjo32,s2022120859,ybkim85\}@cau.ac.kr} \\
}

\begin{document}
\maketitle
\begin{abstract}
Text data augmentation is a complex problem due to the discrete nature of sentences. Although rule-based augmentation methods are widely adopted in real-world applications because of their simplicity, they suffer from potential semantic damage. Previous researchers have suggested easy data augmentation with soft labels (softEDA), employing label smoothing to mitigate this problem. However, finding the best factor for each model and dataset is challenging; therefore, using softEDA in real-world applications is still difficult. In this paper, we propose adapting AutoAugment to solve this problem. The experimental results suggest that the proposed method can boost existing augmentation methods and that rule-based methods can enhance cutting-edge pre-trained language models.  We offer the source code.\footnote{\url{https://github.com/c-juhwan/soft-text-autoaugment}}
\end{abstract}

\section{Introduction}

Data augmentation is a regularization strategy that improves model performance expanding the data held in various ways \cite{hernandez2018data}. In the natural language processing (NLP) field, data augmentation is used in various fields to alleviate data shortages, and various augmentation methods have been proposed accordingly \cite{feng2021survey, li2022data}. For example, image data can be augmented by applying simple rules, such as flipping and rotation, to image data \cite{yang2022image}, and text data can also be augmented, by simple rules such as replacing synonyms and changing the order between words \cite{zhang2015character, wei2019eda}. In addition, a method for augmenting data by generating new text using various deep learning models has also been proposed \cite{sennrich2016improving, wu2019conditional, anaby2020not, yoo2021gpt3mix, zhou2022flipda, dai2023auggpt}. 

However, as these methods often demand training data for fine-tuning before augmentation \cite{zhang2022treemix, li2022data}, it may be challenging to apply them in a low-resource environment \cite{hu2019learning, bayer2022survey, kim2021linda}. Rule-based text data augmentation methods are less costly and easy to implement; thus, they are often used in real-world problems. Despite that, the previously proposed rule-based text data augmentation methods risk not maintaining semantic consistency with original data, which is different from image data \cite{zhao2022epida}, leading to performance degradation.
To relieve this problem, methods that perform data augmentation only through random insertion of punctuation marks have also been proposed \cite{karimi2021aeda}, but they introduce fewer variations compared to easy data augmentation (EDA).  Recently, softEDA \cite{choi2023softeda}, a method applying label smoothing \cite{szegedy2016rethinking} to the augmented data, was proposed to alleviate these drawbacks.

In softEDA, a heuristic grid search was performed for the label smoothing factor (a hyperparameter for performing label smoothing). However, the method based on a heuristic search has the following disadvantages. First, a heuristic search is expensive to execute \cite{bergstra2012random}. Second, although we found the best factor value of the grid, it may not be the global optimum. There could be a better value outside the heuristic search grid; thus, revealing the possible performance gain is difficult.

This paper proposes a method to apply AutoAugment \cite{cubuk2019autoaugment}, a technique to determine the optimal factors in the data augmentation process to alleviate the limitations of previous softEDA methods. By optimizing various arguments of softEDA, it is shown that stable and effective performance improvement is possible compared to the existing rule-based strategy with static factors. In addition, the existing softEDA experiment was conducted on an entire dataset. However, more severe overfitting occurs when the given training data are insufficient \cite{althnian2021impact}, and the scope of performance improvement is greater when additional training data are obtained from a small dataset \cite{prusa2015effect, okimura2022impact}, so data augmentation becomes increasingly crucial in this low-resource environment. Therefore, this study evaluates the proposed method under a low-resource scenario and demonstrates that the proposed method is effective even under data-scarce conditions. In addition, some existing studies have argued that simple rule-based augmentation strategies are less effective in improving the performance of pre-trained language models (PLMs) \cite{longpre2020effective, zhang2022treemix, pluvsvcec2023data}. In this study, we show that through argument optimization, it is possible to improve the performance of not only BERT \cite{devlin2019bert}, the standard PLM, but also DeBERTaV3 \cite{he2022debertav3}, a cutting-edge PLM, through rule-based data augmentation.

\section{Related Work}

Data augmentation of text is primarily performed by augmenting data according to predetermined rules \cite{zhang2015character, belinkov2018synthetic, wei2019eda, karimi2021aeda, choi2023softeda} or using various deep learning models \cite{sennrich2016improving, wu2019conditional, anaby2020not, yoo2021gpt3mix, zhou2022flipda, dai2023auggpt}.  Rule-based data augmentation methods generate new data by performing perturbation in various ways, such as replacing some of the words in a given sentence with synonyms \cite{zhang2015character} or inserting typos at the character  level \cite{belinkov2018synthetic}. The easy data augmentation (EDA) \cite{wei2019eda} technique is a representative rule-based data augmentation method consisting of synonym replacement (SR), random insertion (RI), random swap (RS), and random deletion (RD). However, because such random changes can lead to the loss of semantic consistency, the “an easier data augmentation” (AEDA) technique \cite{karimi2021aeda} consisting only of the RI of six punctuation marks has also been proposed. The softEDA \cite{choi2023softeda} method compensates for the semantic damage caused by EDA by applying label smoothing to the augmented data.

Model-based augmentation methods employ  deep learning models to generate new data. Back-translation \cite{sennrich2016improving} is one of the early model-based methods. It first translates the given data into another language and back-translates it to the original language, generating different expressions with the same concept. Methods based on PLM have also been proposed, and C-BERT \cite{wu2019conditional}, LAMBADA \cite{anaby2020not}, and FlipDA \cite{zhou2022flipda} generate new data using BERT \cite{devlin2019bert}, GPT-2 \cite{radford2019language}, and T5 \cite{raffel2020exploring}, respectively. In addition, GPT-3 \cite{brown2020language} and ChatGPT, which are larger than these PLMs, have been proposed to generate new data \cite{yoo2021gpt3mix, dai2023auggpt}. Other researchers have introduced Mixup \cite{zhang2018mixup} strategy to the NLP field to augment text data \cite{guo2020nonlinear, sun-etal-2020-mixup, yoon2021ssmix}.

Moreover, some previous approaches have tried to apply AutoAugment for NLP. Text AutoAugment \cite{ren2021text}, the work closest to the proposed method, suggested applying AutoAugment to optimize hyperparameters for data augmentation. In addition, DND \cite{kim2021makes} incorporated various data augmentation methods and suggested optimizing two reward terms regarding the difficulty and consistency with the original data. While the proposed work uses AutoAugment to optimize augmentation hyperparameters, we also focus on optimizing label smoothing values for the original and augmented data.

\section{Method}

\subsection{Preliminaries}

The EDA \cite{wei2019eda} method comprises four aforementioned suboperations: SR, RI, RS, and RD. First, SR randomly selects several words in a given sentence and changes them into their synonyms. Second, RI selects a random word in the sentence and inserts its synonym at a random position in the sentence. Third, RS operation randomly selects two words in the sentence and changes their positions. Finally, RD removes each word from the sentence with a predefined probability.

Through these four suboperations, EDA introduces noise to the original data and generates augmented data. Each suboperation has a magnitude of perturbation. For instance, in the case of SR, a higher magnitude leads to the additional replacement of the original words with their synonyms. For each observed data pair $(\bf{x}, \bf{y})$ in the original dataset $\mathcal{D}$, where $\bf{x}$ denotes an input sentence and $\bf{y}$ represents the corresponding label value, the process of EDA can be formulated as follows:

\begin{equation}
\hat{\bf{x}} = \texttt{EDA}(\textbf{x}, p_{\texttt{EDA}}) = 
\newline
\begin{cases} 
\texttt{SR}(\textbf{x}, \alpha_{\texttt{SR}}) \\
\texttt{RI}(\textbf{x}, \alpha_{\texttt{RI}}) \\
\texttt{RS}(\textbf{x}, \alpha_{\texttt{RS}}) \\
\texttt{RD}(\textbf{x}, \alpha_{\texttt{RD}}) \\
\end{cases}
\end{equation}

where $\{\alpha_{\texttt{SR}},\alpha_{\texttt{RI}},\alpha_{\texttt{RS}},\alpha_\texttt{{RD}}\}$ denotes the magnitude of each suboperation, and $p_{\texttt{EDA}} = \{p_{\texttt{SR}}, p_{\texttt{RI}}, p_{\texttt{RS}}, p_{\texttt{RD}}\}$ represents the probability distribution of each suboperation to be selected, which are equal and sum to one. As indicated, EDA only modifies $\bf{x}$, and the label of augmented data is the same as for $\bf{y}$.

The softEDA \cite{choi2023softeda} is a technique that incorporates noise into the label of augmented data through label smoothing \cite{szegedy2016rethinking}. While softEDA follows the previous EDA to augment $\hat{\textbf{x}}$, the following equation defines the process of softEDA, generating a label for augmented data $\hat{\textbf{y}}$:

\begin{equation}
\begin{aligned}
    \hat{\textbf{y}} & = (1-\epsilon_{\textit{aug}})\textbf{y} + \frac{\epsilon_{\textit{aug}}}{N_{\textit{Class}}} \\
    & = 
    \begin{cases}
        (1-\epsilon_{\textit{aug}}) + \frac{\epsilon_{\textit{aug}}}{N_{\textit{Class}}} & \text{if } y = y_i \\
        \frac{\epsilon_{\textit{aug}}}{N_{\textit{Class}}} & \text{Otherwise}
    \end{cases}
\end{aligned}
\end{equation}

where $\epsilon_{\textit{aug}}$ is a smoothing factor for label smoothing.

\subsection{Proposed Method}

Previous EDA and softEDA have numerous augmentation hyperparameters and were primarily fixed or heuristically searched. This paper proposes a method to optimize these hyperparameters by adapting AutoAugment. First, we defined an augmentation policy $\mathcal{P}$ with various factors:

\begin{equation}
\begin{aligned}
\mathcal{P} = \{
p_{\textit{aug}}, p_{\texttt{SR}}, p_{\texttt{RI}}, p_{\texttt{RS}}, p_{\texttt{RD}}, \\
\alpha_{\texttt{SR}},\alpha_{\texttt{RI}},\alpha_{\texttt{RS}},\alpha_{\texttt{RD}}, \\
N_{\textit{aug}}, \epsilon_{\textit{ori}}, \epsilon_{\textit{aug}}
\}
\end{aligned}
\end{equation}

where $p_{\textit{aug}}$ indicates the probability of augmentation, $N_{\textit{aug}}$ refers to the amount of augmented data per original data point, $\epsilon_{\textit{ori}}$ represents a label smoothing factor for the original data, different from $\epsilon_{\textit{aug}}$. Following Text AutoAugment \cite{ren2021text}, we optimized the proposed policy based on sequential model-based global optimization \cite{bergstra2011algorithms}. Finding the optimal augmentation parameter for each model and dataset through this adaptation of AutoAugment with softEDA is more beneficial than inefficient grid search.

\section{Experiment}

\begin{table*}[]
\resizebox{0.98\textwidth}{!}{ 

\begin{tabular}{l|cccccccc}

\Xhline{3\arrayrulewidth}
                  & \textbf{SST2}                                                                    & \textbf{SST5}                                                                    & \textbf{CoLA}                                                                    & \textbf{SUBJ}                                                                    & \textbf{TREC}                                                                    & \textbf{MR}                                                                      & \textbf{CR}                                                                      & \textbf{PC}                                                                      \\ \hline\hline
\textbf{BERT} w/o Aug      & \begin{tabular}[c]{@{}c@{}}$80.46_{1.84}$\\ $86.08_{1.03}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$35.13_{0.74}$\\ $43.64_{0.50}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$71.49_{1.40}$\\ $75.50_{0.58}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$92.85_{0.44}$\\ $95.07_{\bm{0.22}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$78.42_{1.30}$\\ $93.27_{0.42}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$72.11_{1.39}$\\ $81.29_{0.52}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$79.88_{0.82}$\\ $87.53_{0.60}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$88.12_{0.58}$\\ $91.15_{0.21}$\end{tabular} \\ \hline
w/ EDA            & \begin{tabular}[c]{@{}c@{}}$80.76_{1.39}$\\ $86.71_{0.63}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$36.63_{1.33}$\\ $45.08_{1.16}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$\textcolor{gray}{70.70}_{0.98}$\\ $\textcolor{gray}{73.18}_{0.52}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$93.39_{\bm{0.25}}$\\ $\textcolor{gray}{94.69}_{0.33}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$81.56_{1.71}$\\ $93.99_{1.05}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$73.18_{1.36}$\\ $\textcolor{gray}{80.41}_{0.29}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$\textcolor{gray}{79.54}_{1.15}$\\ $87.71_{0.57}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$89.64_{0.80}$\\ $\textcolor{gray}{90.81}_{0.40}$\end{tabular} \\ \hline
w/ AEDA           & \begin{tabular}[c]{@{}c@{}}$80.96_{1.63}$\\ $86.66_{0.63}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$36.54_{0.97}$\\ $44.53_{1.02}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$72.24_{1.85}$\\ $\textcolor{gray}{74.44}_{0.41}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$93.29_{0.23}$\\ $\textcolor{gray}{94.60}_{0.48}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$81.27_{2.19}$\\ $93.87_{0.75}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$74.37_{2.84}$\\ $81.57_{\bm{0.15}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$80.67_{1.64}$\\ $87.66_{0.55}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$88.75_{0.90}$\\ $\textcolor{gray}{91.03}_{0.31}$\end{tabular} \\ \hline
w/ softEDA        & \begin{tabular}[c]{@{}c@{}}$80.80_{3.22}$\\ $87.84_{0.65}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$37.13_{1.60}$\\ $45.04_{1.28}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$72.41_{0.95}$\\ $\textcolor{gray}{74.16}_{0.99}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$93.24_{0.40}$\\ $\textcolor{gray}{94.85}_{0.39}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$82.92_{1.70}$\\ $94.68_{0.51}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$74.40_{1.27}$\\ $\textcolor{gray}{81.16}_{0.88}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$\textcolor{gray}{78.95}_{2.65}$\\ $87.94_{0.85}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$88.82_{1.63}$\\ $\textcolor{gray}{91.12}_{0.63}$\end{tabular} \\ \hline
w/ Ours           & \begin{tabular}[c]{@{}c@{}}${\bm{85.48}}_{0.57}$\\ ${\bm{88.53}}_{\bm{0.27}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{39.88}}_{0.41}$\\ ${\bm{46.16}}_{0.63}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{74.63}}_{\bm{0.33}}$\\ ${\bm{76.66}}_{0.81}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{94.10}}_{0.35}$\\ ${\bm{95.54}}_{0.33}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{85.88}}_{\bm{1.06}}$\\ ${\bm{95.17}}_{0.54}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{79.32}}_{\bm{0.37}}$\\ ${\bm{83.10}}_{0.34}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{86.49}}_{\bm{0.22}}$\\ ${\bm{89.98}}_{\bm{0.25}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{91.54}}_{\bm{0.11}}$\\ ${\bm{92.16}}_{0.19}$\end{tabular} \\ \hline
w/ Ours w/o LS    & \begin{tabular}[c]{@{}c@{}}$84.71_{\bm{0.44}}$\\ $88.13_{0.48}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$39.22_{\bm{0.38}}$\\ $45.45_{\bm{0.39}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$73.80_{0.79}$\\ $76.30_{\bm{0.34}}$\end{tabular}   & \begin{tabular}[c]{@{}c@{}}$93.71_{0.35}$\\ $95.15_{\bm{0.22}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$84.85_{1.40}$\\ $94.70_{\bm{0.46}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$77.86_{0.53}$\\ $82.19_{0.60}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$85.70_{0.88}$\\ $89.66_{0.35}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$91.13_{0.19}$\\ $91.98_{\bm{0.18}}$\end{tabular} \\ \hline\hline

\textbf{DeBERTaV3} w/o Aug & \begin{tabular}[c]{@{}c@{}}$88.36_{0.36}$\\ $92.59_{0.73}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$35.95_{1.69}$\\ $48.77_{1.52}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$72.62_{4.24}$\\ $82.21_{0.82}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$92.23_{0.24}$\\ $94.66_{0.22}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$80.19_{3.23}$\\ $94.06_{0.43}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$82.84_{0.39}$\\ $86.22_{0.37}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$85.61_{1.20}$\\ $91.40_{0.36}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$91.22_{0.43}$\\ $91.85_{0.26}$\end{tabular} \\ \hline

w/ EDA            & \begin{tabular}[c]{@{}c@{}}$\textcolor{gray}{86.61}_{0.70}$\\ $93.25_{0.55}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$37.64_{1.23}$\\ $49.04_{0.78}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$74.83_{1.10}$\\ $\textcolor{gray}{79.24}_{0.66}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$92.85_{0.48}$\\ $94.81_{0.53}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$83.65_{1.84}$\\ $94.33_{0.99}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$83.18_{\bm{0.32}}$\\ $86.71_{0.65}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$\textcolor{gray}{84.86}_{0.73}$\\ $\textcolor{gray}{91.24}_{0.39}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$\textcolor{gray}{90.51}_{0.47}$\\ $92.3_{0.15}$\end{tabular}  \\ \hline

w/ AEDA           & \begin{tabular}[c]{@{}c@{}}$88.44_{0.80}$\\ $\textcolor{gray}{92.54}_{0.78}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$36.87_{2.88}$\\ $49.16_{0.83}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$79.29_{0.65}$\\ $82.78_{\bm{0.40}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$92.81_{0.47}$\\ $94.92_{0.58}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$84.17_{0.79}$\\ $94.45_{0.80}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$82.87_{0.75}$\\ $\textcolor{gray}{85.77}_{1.63}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$85.76_{1.37}$\\ $\textcolor{gray}{91.09}_{0.49}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$90.61_{0.49}$\\ $92.29_{\bm{0.11}}$\end{tabular} \\ \hline

w/ softEDA        & \begin{tabular}[c]{@{}c@{}}$88.94_{1.03}$\\ $93.12_{1.05}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$38.37_{1.65}$\\ $50.34_{1.44}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$79.40_{1.51}$\\ $\textcolor{gray}{78.97}_{1.16}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$92.90_{1.08}$\\ $94.77_{0.21}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$84.58_{1.29}$\\ $94.71_{0.69}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$83.50_{0.65}$\\ $87.02_{0.50}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$86.33_{1.65}$\\ $91.81_{0.76}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$91.28_{0.82}$\\ $92.16_{0.20}$\end{tabular} \\ \hline

w/ Ours           & \begin{tabular}[c]{@{}c@{}}${\bm{91.38}}_{0.32}$\\ ${\bm{93.94}}_{\bm{0.30}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{42.92}}_{0.52}$\\ ${\bm{52.77}}_{\bm{0.62}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{82.56}}_{0.51}$\\ ${\bm{84.32}}_{0.49}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{94.47}}_{0.26}$\\ ${\bm{95.29}}_{0.31}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{87.70}}_{0.90}$\\ ${\bm{94.92}}_{0.62}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{85.31}}_{0.79}$\\ ${\bm{87.96}}_{\bm{0.17}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{89.95}}_{\bm{0.51}}$\\ ${\bm{92.46}}_{\bm{0.18}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}${\bm{92.32}}_{\bm{0.19}}$\\ ${\bm{92.72}}_{0.40}$\end{tabular} \\ \hline

w/ Ours w/o LS    & \begin{tabular}[c]{@{}c@{}}$90.47_{\bm{0.26}}$\\ $93.40_{0.58}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$42.44_{\bm{0.49}}$\\ $52.54_{0.66}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$82.10_{\bm{0.43}}$\\ $83.67_{0.86}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$94.22_{\bm{0.15}}$\\ $95.15_{\bm{0.12}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$86.57_{\bm{0.61}}$\\ $94.92_{\bm{0.18}}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$85.07_{0.58}$\\ $87.41_{0.37}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$89.47_{0.67}$\\ $92.28_{0.27}$\end{tabular} & \begin{tabular}[c]{@{}c@{}}$92.22_{0.21}$\\ $92.49_{0.33}$\end{tabular} \\ \Xhline{3\arrayrulewidth}

\end{tabular}
}
\caption{Experimental results. Each experiment has been repeated five times and the statistics are presented in $\textit{mean}_{\textit{std}}$ format. The upper side of each column denotes the results when $N_{\textit{Train}}=100$, and the lower side shows the results when $N_{\textit{Train}}=500$. The best mean and standard deviation values for each model and dataset are boldfaced. Results that reported a lower mean value than the baseline are gray.}
\label{tab-main}
\end{table*}

\subsection{Datasets and Low-resource Setting}

Eight text classification datasets were used to evaluate the proposed method. The SST2, SST5 \cite{socher2013recursive} and MR \cite{pang-etal-2002-thumbs} sentiment classification tasks are from movie reviews. The CoLA \cite{warstadt2019neural} binary classification dataset measures the linguistic acceptability of a given sentence. The SUBJ \cite{pang2004sentimental} binary classification dataset deals with the subjectivity of a sentence. PC \cite{ganapathibhotla-liu-2008-mining}, and CR \cite{hu2004mining, qian2015automated} are datasets constructed from customer reviews. In addition, the TREC \cite{li2002learning} multiclass text classification dataset is about the question type of given text. Dataset specifications can be found in Appendix~\ref{sec:appendix-dataset}.

Data augmentation becomes more important when the given data is deficient than when sufficient data can be accessed \cite{chen2023empirical}. To simulate a more challenging scenario, we evaluated the proposed method with only 100 and 500 randomly selected original data from each dataset.

\subsection{Baselines}

To validate the claim that hyperparameter optimization for the augmentation method is effective in enhancing model performance, we compared our approach with previous rule-based data augmentation methods with fixed hyperparameters. We compared the proposed method against the previous EDA, AEDA, and softEDA methods with fixed hyperparameters.

Recent studies suggest that simple rule-based augmentation methods are insufficient to enhance PLM-based models \cite{longpre2020effective, zhang2022treemix, pluvsvcec2023data}. In addition, validating the newly proposed augmentation method using cutting-edge models, not just models like BERT, is necessary \cite{zhou2022flipda}. Therefore, we adopted BERT and DeBERTaV3 \cite{he2022debertav3}, an improvement of DeBERTa \cite{he2020deberta} as the baseline model for evaluation.

\subsection{Main Results}

Table~\ref{tab-main} reports the experimental results. Previously proposed augmentation methods have faced marginal gain, or even performance degradation. Especially, softEDA has a high standard deviation compared to other methods, indicating that softEDA has difficulty being effective within a single fixed hyperparameter and requires optimization for hyperparameters. Whereas, the proposed method exhibits a stable and remarkable performance improvement within every setting, including those where other methods had performance degradation or marginal gains. This finding suggests enhancing extensive and cutting-edge PLMs with simple augmentation methods is achievable under the carefully designed data augmentation policy and hyperparameter optimization strategy. Furthermore, it is shown that our strategy has remarkably low standard deviation values compared to other techniques, showcasing that our approach is robust against statistical differences and valuable for practical application in low-resource text classification problems.

\subsection{Ablation Study}

One may wonder whether the performance improvement reported in Table~\ref{tab-main} is solely caused by the adaptation of AutoAugment, rather than the label smoothing of softEDA. To validate the effectiveness of label smoothing, we conducted an ablation study where label smoothing is not applied (i.e., $\epsilon_{\textit{ori}} = \epsilon_{\textit{aug}} = 0$). This setting is equal to optimizing only factors of EDA. ``w/ Ours w/o LS'' row of Table~\ref{tab-main} presents the experimental results, revealing that the proposed method without label smoothing is less effective than the proposed method. This finding supports that the label smoothing optimization introduced by softEDA plays a crucial role in enhancing the model.

\section{Conclusion}

This paper proposed a method to optimize various hyperparameters of rule-based text augmentation methods. The experimental results suggest that the proposed method is effective and stable, and that rule-based augmentation methods can improve cutting-edge PLMs with proper hyperparameter optimization. Future work may extend this approach to other tasks, such as natural language inference, which is more complex than the single-sentence classification conducted in this paper.

\section*{Limitations}
This paper used AutoAugment to optimize the rule-based data augmentation method. The primary weakness of AutoAugment is the computational overhead from the searching process \cite{zhang2022good}. However, under low-resource situations, where the necessity of data augmentation is emphasized, this problem can be diminished as the time consumption of the search process decreases.

\section*{Ethics Statement}
This paper proposes an optimized rule-based augmentation method. These rule-based methods are more ethically stable than model-based approaches, as the modification is performed under predefined rules. For example, back-translation can be easily exposed to the potential bias of the translation model. Methods based on PLMs also share this concern. However, rule-based augmentation methods, including the proposed method, perform modifications within a given sentence and are less likely to be exposed to unintentional bias.

\section*{Acknowledgements}
This research was supported by Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education(NRF-2022R1C1C1008534), and Institute for Information \& communications Technology Planning \& Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program, Chung-Ang University).

\bibliography{custom}

\appendix

\section{Dataset Specification}
\label{sec:appendix-dataset}

\begin{table}[h!]
\centering
\resizebox{0.8\columnwidth}{!}{
\begin{tabular}{c|ccc}
\Xhline{3\arrayrulewidth}
Dataset & $N_{\textit{Class}}$ & $N_{\textit{Train}}$ & $N_{\textit{Test}}$ \\ \hline\hline
SST2    & 2      & 6.9K   & 1.8K  \\
SST5    & 5      & 8.5K   & 2.2K  \\
CoLA    & 2      & 8.5K   & 0.5K  \\
SUBJ    & 6      & 8K     & 2K    \\
TREC    & 2      & 5.5K   & 0.5K  \\
MR      & 2      & 9.5K   & 1.1K  \\
CR      & 2      & 3.0K   & 0.8K  \\
PC      & 2      & 39K    & 4.5K  \\\Xhline{2\arrayrulewidth}
\end{tabular}
}
\caption{Specification of each dataset used for the experiment.}
\label{tab-dataset}
\end{table}

\begin{table*}[ht!]
\centering
\resizebox{0.8\textwidth}{!}{
\begin{tabular}{l|lllllll}
\Xhline{3\arrayrulewidth}
                & \multicolumn{1}{c}{SST2}     & \multicolumn{1}{c}{CR}       & \multicolumn{1}{c}{MR}       & \multicolumn{1}{c}{TREC}     & \multicolumn{1}{c}{SUBJ}     & \multicolumn{1}{c}{PC}       & \multicolumn{1}{c}{CoLA}     \\ \hline\hline
BERT w/o Aug    & 89.74                        & 89.08                        & 84.28                        & 95.47                        & 96.18                        & 93.44                        & 75.38                        \\
w/ EDA          & +0.71                        & {\color[HTML]{9B9B9B} -0.41} & {\color[HTML]{9B9B9B} -0.92} & +0.51                        & {\color[HTML]{9B9B9B} -0.35} & +0.58                        & {\color[HTML]{9B9B9B} -0.45} \\
w/ AEDA         & +0.22                        & +1.84                        & \textbf{+0.19}               & {\color[HTML]{9B9B9B} -0.67} & {\color[HTML]{9B9B9B} -0.30} & {\color[HTML]{9B9B9B} -0.15} & {\color[HTML]{9B9B9B} -0.34} \\
w/ softEDA 0.1  & {\color[HTML]{9B9B9B} -0.11} & +0.29                        & {\color[HTML]{9B9B9B} -1.10} & {\color[HTML]{9B9B9B} -1.45} & \textbf{+0.15}               & +0.43                        & +1.34                        \\
w/ softEDA 0.15 & {\color[HTML]{9B9B9B} -0.22} & +0.66                        & {\color[HTML]{9B9B9B} -0.46} & {\color[HTML]{9B9B9B} -0.47} & {\color[HTML]{9B9B9B} -0.50} & {\color[HTML]{9B9B9B} -0.01} & +0.02                        \\
w/ softEDA 0.2  & {\color[HTML]{9B9B9B} -0.12} & \textbf{+2.10}               & \textbf{+0.19}               & {\color[HTML]{9B9B9B} -0.27} & +0.05                        & +0.43                        & +0.81                        \\
w/ softEDA 0.25 & {\color[HTML]{9B9B9B} -0.23} & \textbf{+2.10}               & {\color[HTML]{9B9B9B} -0.92} & \textbf{+1.17}               & {\color[HTML]{9B9B9B} -0.10} & \textbf{+0.67}               & \textbf{+1.50}               \\
w/ softEDA 0.3  & \textbf{+0.83}               & {\color[HTML]{9B9B9B} -0.90} & {\color[HTML]{9B9B9B} -1.80} & {\color[HTML]{9B9B9B} -0.78} & +0.00                        & \textbf{+0.67}               & +0.23                        \\ \Xhline{3\arrayrulewidth}
\end{tabular}}
\caption{Results of softEDA for the BERT model reported in the softEDA paper. The best scores for each dataset are boldfaced. Scores lower than the baseline are gray.}
\label{tab-softeda}
\end{table*}

\section{Implementation Details}
\label{sec:appendix-detail}

We used PyTorch \cite{paszke2019pytorch} and Huggingface Transformers \cite{wolf2020transformers} to implement the model and evaluation process. We used \texttt{bert-base-cased} and \texttt{microsoft/deberta-v3-base} for the BERT and DeBERTaV3 models. Every model was trained using the Adam optimizer with a batch size of 32 and a learning rate of 5e-5 for ten epochs, with early stopping with a patience value of 5, conditioned on best validation accuracy. The training procedure was performed on a single Nvidia RTX 3090 GPU.

For the baseline method implementation, we used TextAugment library \cite{marivate2020improving} for EDA, and softEDA was built on it. The library did not have an implementation for AEDA; thus, we implemented it separately. We used ray tune \cite{liaw2018tune} to implement the proposed method. Please refer to the attached code for more information.

\section{Analysis of softEDA}
\label{sec:appendix-analysis}

We investigated the experimental results of the softEDA paper. Table~\ref{tab-softeda} presents the experimental results reported in the appendix of the softEDA paper. The results suggest that, although softEDA can potentially enhance model performance, it is problematic to determine the optimal label smoothing factor for each model and dataset. Performance degradation compared to the baseline was also observed where the factor is improper for each setup. This finding motivated us to determine a better solution for finding optimal factors than a heuristic search. Furthermore, the authors performed the experiment on the full dataset. In contrast, we conducted the experiment through low-resource scenarios, which is more challenging for model.

\end{document}